# Can language models boost the power of randomized experiments without statistical bias?

### v1

@@ 5.4 Insights on CALM
In the previous section, the CALM method demonstrated efficiency gains over the benchmark methods. Here, we provide further insights into CALM by addressing two key questions: How informative are LLM-based predictions through $\omega_{t}(x)$, and how informative are unstructured covariates in Section 5.4.1 and Section 5.4.2, respectively.
@@ 5.4.1 How informative are LLM-based predictions through $\omega_{t}(x)$?
To start, we provide some insights on the informativeness of LLM-based predictions through the lens of the calibration weights $\omega_{t}(x)$. Specifically, we examine the relationship between the LLM-predicted outcomes and the observed outcomes across different covariate strata. This analysis is motivated by the inherent structure of the CALM method, which selectively borrows information from distinct regions of the covariate space, depending on where the LLM predictions are most informative.
@@
5.4.2 How informative are unstructured covariates?
As outlined in the motivation for our proposed method, leveraging LLMs to generate outcome predictions offers two main advantages: (i) the seamless integration of unstructured covariates and (ii) the ability to exploit the knowledge embedded in pre-trained LLMs. These considerations naturally raise two questions: To what extent do unstructured covariates enhance the informativeness of LLM predictions, and consequently improve the efficiency of treatment effect estimation? And how much of the gain can be attributed to the prior knowledge encoded in pre-trained LLMs themselves?
Figure 5 demonstrates that zero-shot learning—whether or not unstructured covariates are included in the prompt—yields higher estimation efficiency compared to using standard machine learning methods alone. Moreover, incorporating unstructured covariates into the zero-shot prompt provides an additional efficiency gain. These results suggest that the pre-trained LLM (gpt-4o-mini in this setting) encodes knowledge that improves the efficiency of ATE estimation. Even without informative unstructured inputs, the LLM’s exposure to large and diverse training corpora may enable it to capture latent associations between covariates and outcomes, which CALM can then exploit to enhance efficiency. Incorporating unstructured covariates into the prompt further amplifies this benefit.
@@ 7 Conclusion
In this manuscript, we introduced CALM, a statistically principled framework for leveraging LLM-generated prognostic predictions in randomized experiments. By residualizing and heterogeneously calibrating LLM predictions, CALM yields consistent estimators with valid inference even when zero- and few-shot predictions are biased, and it empirically improves efficiency relative to covariate-adjusted AIPW for mean potential outcomes; extensions to average and heterogeneous treatment effects follow directly. In simulations calibrated to the BRIGHTEN RCT, CALM reduces variance and increases power for detecting treatment-effect heterogeneity while remaining robust to prompt design.

### v2

@@ 5.4 Insights on CALM
In the previous section, the CALM method demonstrated efficiency gains over the benchmark methods. Here, we provide further insights into CALM by addressing two key questions: How informative are pre-trained LLMs, and how informative are LLM-based predictions through $\omega_{t}(x)$ in Section 5.4.1 and Section 5.4.2 , respectively.
5.4.1 Does pre-trained LLM contain relevant information?
As outlined in the motivation for our proposed method, leveraging LLMs to generate outcome predictions offers two main advantages: (i) the ability to exploit the knowledge embedded in pre-trained LLMs, and (ii) the seamless integration of unstructured covariates. These considerations naturally raise two questions: How much of the gain can be attributed to the prior knowledge encoded in pre-trained LLMs themselves? And to what extent do unstructured covariates enhance the informativeness of LLM predictions, and consequently improve the efficiency of treatment effect estimation?
Figure 4 demonstrates that our proposed method with zero-shot learning–whether or not unstructured covariates are included in the prompt–yields higher estimation efficiency compared to using standard machine learning methods alone. Moreover, incorporating unstructured covariates into the zero-shot prompt provides an additional efficiency gain. These results suggest that the pre-trained LLM (gpt-4o-mini in this setting) encodes knowledge that improves the efficiency of ATE estimation. Even without informative unstructured inputs, the LLM’s exposure to large and diverse training corpora may enable it to capture latent associations between covariates and outcomes, which CALM can then exploit to enhance efficiency. Incorporating unstructured covariates into the prompt further amplifies this benefit.
@@ 5.4.2 How informative are LLM-based predictions through $\omega_{t}(x)$?
Next, we provide some insights on the informativeness of LLM-based predictions through the lens of the calibration weights $\omega_{t}(x)$. Specifically, we examine the relationship between the LLM-predicted outcomes and the observed outcomes across different covariate strata. This analysis is motivated by the inherent structure of the CALM method, which selectively borrows information from distinct regions of the covariate space, depending on where the LLM predictions are most informative.
@@ 7 Conclusion
In this manuscript, we introduced CALM, a statistically principled framework for leveraging LLM-generated predictions in randomized experiments. By residualizing and heterogeneously calibrating LLM predictions, CALM yields consistent estimators with valid inference even when zero- and few-shot predictions are biased, and it empirically improves efficiency relative to covariate-adjusted AIPW for mean potential outcomes; extensions to average and heterogeneous treatment effects follow directly. In simulations calibrated to the BRIGHTEN RCT, CALM reduces variance and increases power for detecting treatment-effect heterogeneity while remaining robust to prompt design.
Data availability
The BRIGHTEN study data are publicly accessible upon request via Synapse:
https://www.synapse.org/Synapse:syn10848316/wiki/548727 (Synapse Project:syn10848316).